# E-values, Multiple Testing and Beyond

### v1

@@ 4.2 Fair multiple testing
$T_{g}=\sup\left\{0<t<0.5\colon\frac{1+\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{p_{i}\geq 1-t\}}{1\vee\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{p_{i}\leq t\}}\leq\alpha\right\}$ (15)
be the rejection threshold for the $g$th group with $1\leq g\leq G$. Define
$e_{i}=\frac{n_{g}w_{i}\mathbbm{1}\{p_{i}\leq T_{g}\}}{1+\sum_{j\in\mathcal{G}_{g}}\mathbbm{1}\{p_{j}\geq 1-T_{g}\}},$ (16)
for $i\in\mathcal{G}_{g}$, where $w_{i}>0$ represents the weight for the $i$th hypothesis, which will be specified in Section 4.2.1. After collecting the e-values from each group, we implement the e-BH procedure at level $\alpha$. The testing procedure is summarized in Algorithm 2.
Algorithm 2 Fair multiple testing
p-values $p_{1},\dots,p_{n}$, group indices $\mathcal{G}_{1},\dots,\mathcal{G}_{G}$, significance level $\alpha$
for $g=1,\dots,G$ do
     Implement the BC procedure utilizing the p-values $\{p_{i}\colon i\in\mathcal{G}_{g}\}$ at the level $\alpha$.
     Calculate the threshold $T_{g}$ using (15).
     for $i\in\mathcal{G}_{g}$ do
         Calculate the e-value $e_{i}$ using (16).
     end for
end for
Assemble the e-values from all groups.
Run the e-BH procedure utilizing the assembled e-values at the level $\alpha$.
The indices of rejected hypotheses.
4.2.1 Choice of weights and FDR control
Controlling the overall FDR requires ensuring that the e-values defined in (16) satisfy Condition (3). One strategy is to set $w_{i}=1$ for all $i$. Another strategy takes into account the group size by setting $w_{i}=n/(Gn_{g})$ for all $i\in\mathcal{G}_{g}$ and $g=1,\dots,G$, where $n_{g}=|\mathcal{G}_{g}|$ is the cardinality of the $g$th group. It can be verified that Condition (3) is satisfied by both strategies. The e-BH procedures based on these two weight choices are referred to as eBH_1 and eBH_2, respectively.
Theorem 6.
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and satisfy Condition (8), and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. Further suppose the weights $w_{i}$ are independent of the p-values and satisfy
$\sum_{g=1}^{G}n_{g}\max_{i\in\mathcal{G}_{g}}w_{i}\leq n.$ (17)
Then, the e-values specified in (16) satisfy Condition (3). Hence, the corresponding e-BH procedure controls the FDR in finite sample.
Our simulations show that eBH_1 and eBH_2 frequently suffer from low statistical power. To enhance efficiency, we suggest utilizing a data-dependent weight selection approach inspired by the leave-one-out technique (Barber et al., 2020).
Denote the p-values in the $g$th group by $\mathbf{p}_{g}=\{p_{i}\}_{i\in\mathcal{G}_{g}}$. Write $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$, and let $\mathbf{p}_{g,i}$ for $i\in\mathcal{G}_{g}$ be the collection of p-values obtained by replacing $p_{i}$ with $\tilde{p}_{i}$ in $\mathbf{p}_{g}$. By viewing $T_{g}$ as a functions of $\mathbf{p}_{g}$, we define $T_{g,i}=T_{g}(\mathbf{p}_{g,i})$, i.e., the threshold of the BC procedure applied to the set of p-values $\mathbf{p}_{g,i}$. We define the data-dependent weights as
$w_{i}=\frac{\frac{n}{n_{g}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{p_{j}\geq 1-T_{g}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{p_{j}\geq 1-T_{g}\}\right)+\sum_{g^{\prime}\neq g}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{p_{j}\geq 1-T_{g^{\prime},j}\}},$ (18)
for $i\in\mathcal{G}_{g}$. When $G=2$, we have
for $i\in\mathcal{G}_{1}$ and
for $i\in\mathcal{G}_{2}$. The e-BH procedure based on the weights specified in (18) is referred to as the eBH_Ada method in the following discussions. The theorem below states that eBH_Ada has finite sample FDR control.
Theorem 7.
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and satisfy Condition (8), and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. Then, the e-values specified in (16) with the weights defined via (18) satisfy Condition (3). Hence, the corresponding e-BH procedure controls the FDR in finite sample.
4.2.2 Numerical studies
We shall compare the finite sample performance of the proposed method with two naive approaches through simulations. The first method disregards the group information and directly applies the BC procedure to all p-values. We refer to this method as BC_Com for future reference. BC_Com has two shortcomings. Firstly, it may fail to control the group-wise FDRs, as illustrated in Setting E2. Secondly, it fails to ensure comparable power across different groups, resulting in the possibility of one group having high power while the other group has nearly zero power, as illustrated in Setting E1. The same issue is also encountered by eBH_1. An alternative approach involves implementing the BC procedure for each group separately and combining all rejections. We call this method BC_Sep. Although BC_Sep effectively controls the FDR for individual groups, it does not guarantee the overall FDR control.
Remark 2.
We also tried BC_Sep at the level $\alpha/G$, a naive method that controls both the group-wise and overall FDRs. Indeed, let $\hat{n}_{ga}$ be the number of rejections for the $g$th group, and denote the number of false rejections for the $g$th group by $\hat{n}_{g0}$. Then we have $\mathbb{E}[\hat{n}_{g0}/(1\vee\hat{n}_{ga})]\leq\alpha/G$, which implies that
$\mathbb{E}\left[\frac{\sum_{g=1}^{G}\hat{n}_{g0}}{1\vee\sum_{g=1}^{G}\hat{n}_{ga}}\right]\leq\mathbb{E}\left[\sum_{g=1}^{G}\frac{\hat{n}_{g0}}{1\vee\hat{n}_{ga}}\right]\leq\alpha.$
However, this method has nearly zero power in all our simulation settings. Therefore, we have decided not to include its results in the tables below.
[…]
In all settings, we assume that the p-values follow the uniform distribution on $[0,1]$ under the null. For the first group, the p-value is assumed to follow $\mathrm{Unif}[0,u_{1}]$ under the alternatives, while for the second group, it follows $\mathrm{Unif}[0,u_{2}]$ under the alternatives. The parameter values for different settings are detailed in Table 2.
Setting $n_{1}$ $n_{1a}$ $u_{1}$ $n_{2}$ $n_{2a}$ $u_{2}$
E1 $100$ $20$ $1.5\times 10^{-2}$ $1000$ $20$ $1.5\times 10^{-3}$
E2 $100$ $20$ $1.5\times 10^{-3}$ $1000$ $20$ $1.5\times 10^{-3}$
Table 2: Parameter settings for the case of $G=2$. Here, $n_{g}$ represents the number of hypotheses for the $g$th group; $n_{ga}$ denotes the number of non-null hypotheses in the $g$th group. The parameter $u_{g}$ is the upper bound of the uniform distribution for the p-values under the alternatives for the $g$th group.
Setting E1 corresponds to a scenario in which, for instance, the first group consists of ethnic minorities while the second group comprises of majorities. The number of non-nulls are the same across the two groups. The alternative p-values in the first group are larger than those in the second group on average as $u_{1}>u_{2}$. Table 3 summarizes the results for this setting. BC_Com exhibits high power for the second group, yet its power in the first group are rather low. This is because the non-null p-values from the first group are not sufficiently small, and a combined analysis of the two groups demands a lower threshold and thus fails to reject them. Additionally, BC_Sep has a slightly inflated overall FDR in this case.
[…]
BC_Com 0.144 0.041 0.248 0.023 0.012 0.025
BC_Sep 0.296 0.352 0.24 0.053 0.036 0.025
eBH_1 0.07 0 0.141 0.02 0 0.02
eBH_2 0.078 0.077 0.078 0.008 0.008 0.008
eBH_Ada 0.181 0.212 0.149 0.028 0.021 0.015
Table 3: FDR and power for Setting E1, where the nominal FDR level is 5%.
Setting E2 is the same as Setting E1 except that $u_{1}=u_{2}=1.5\times 10^{-2}$. The results for Setting E2 are presented in Table 4. We observe that BC_Com fails to control the FDR for the second group, which can be explained as follows. Due to the fact that the non-null p-values have a similar scale for both groups and the sample size of the first group is significantly smaller than that of the second group, BC_Com has a higher threshold compared to the BC procedure applied only to the second group. This can result in an FDR inflation in the second group for BC_Com. We also observe that BC_Sep suffers from an overall FDR inflation. In contrast, all three variants of the e-BH procedure control the group-wise and overall FDRs at the desired level. eBH_Ada has a much higher power than the other two e-value based methods.
[…]
BC_Com 0.605 0.605 0.606 0.039 0.006 0.067
BC_Sep 0.573 0.906 0.24 0.057 0.048 0.025
eBH_1 0.07 0 0.141 0.02 0 0.02
eBH_2 0.216 0.217 0.215 0.017 0.011 0.022
eBH_Ada 0.384 0.552 0.216 0.036 0.029 0.023
Table 4: FDR and power for Setting E2, where the nominal FDR level is 5%.
We present the results for $G=4$ in Appendix C.1, where we consider three different scenarios (Settings F1-F3). In particular, we note that BC_Com suffers from a severe FDR inflation with the empirical FDR being 0.312 at the 5% target level in Setting F2. In Setting F3, BC_Sep has an empirical overall FDR being 0.346, which is much higher than the 20% level.
To summarize, as seen in Settings E2 and F2, BC_Com has no guarantee in controlling the group-wise FDR. On the other hand, BC_Sep fails to control the overall FDR, as observed in all the settings, particularly Settings E2 and F3 in Appendix C.1. The e-BH based approaches provide both group-wise and overall FDR control, making them fair according to Definition 1. However, eBH_1 and eBH_2 may suffer from power loss under certain scenarios. In contrast, eBH_Ada demonstrates consistent effectiveness across all settings by achieving FDR fairness and reasonable power.
4.3 Structure adaptive multiple testing
Having access to various types of auxiliary information that reflect the structural relationships among hypotheses is becoming increasingly common. Taking advantage of such auxiliary information can improve the statistical power in multiple testing. In this section, we consider the scenarios where, in addition to the p-value $p_{i}$, there is associated structural information in the form of a covariate $x_{i}$ for each hypothesis. This side information represents heterogeneity among the p-values and may affect the prior probabilities of the null hypotheses being true or the signal strength under alternatives. Our goal is to develop a multiple testing procedure that can incorporate such external structural information to improve statistical power and has guaranteed FDR control in finite sample. The high-level idea behind our approach is to relax the p-value thresholds for hypotheses that are more likely to be non-null and tighten the thresholds for the other hypotheses through the use of hypothesis-specific rejection rule, i.e., $\varphi_{i}(p_{i})\leq t$, so that the FDR can be controlled.
Our proposed method combines the cross-fitting technique (Ignatiadis and Huber, 2021) (a sample-splitting and fitting approach that enables learning the hypothesis-specific rejection function $\varphi_{i}$ without overfitting as long as the hypotheses can be partitioned into independent folds) with the GBC procedure introduced in Section 3. First, we randomly split the data into $G$ distinct groups, denoted as $\{\mathcal{G}_{g}\colon g=1,\dots,G\}$, where $\bigcup_{g=1}^{G}\mathcal{G}_{g}=[n]$ and $\mathcal{G}_{g}\cap\mathcal{G}_{g^{\prime}}=\emptyset$ for $g\neq g^{\prime}$. We estimate the rejection function $\varphi_{i}$ for the hypothesis $i$ in the $g$th group using the data from all other groups, which ensures that the estimated rejection function is independent of the p-values in group $g$. We then apply the GBC procedure based on the estimated rejection functions separately to each group and obtain the corresponding e-values. Finally, we collect all the e-values and apply the e-BH procedure at the target level to control the FDR.
To describe the cross-fitting procedure, let us assume that $\varphi_{i}(p)=\varphi(p,x_{i};\beta)$ for some unknown parameter $\beta$ that needs to be estimated from the data. We define the cross-fitted estimate as
$\hat{\beta}_{-g}=\operatorname*{arg\,min}_{\beta}\sum_{i\notin\mathcal{G}_{g}}\mathcal{L}(p_{i},x_{i},\beta).$
Here $\mathcal{L}$ is some loss function such as negative log-likelihood under the two-group mixture model in Example 3, in which case we assume $\pi_{i}(p)=\pi(p,x_{i};\beta_{1})$, $f_{1,i}(p)=f(p,x_{i};\beta_{2})$ and $\beta=(\beta_{1},\beta_{2}).$ It should be noted that the choice of the functional form of $\varphi_{i}$ is highly flexible and does not impact the FDR control due to cross-fitting that avoids over-fitting. Given $\hat{\beta}_{-g}$, we define
$\hat{\varphi}_{i}(p)=\varphi(p,x_{i};\hat{\beta}_{-g})\quad\text{for}\quad i\in\mathcal{G}_{g}.$
Next, we apply the GBC procedure using the cross-fitted functions $\{\hat{\varphi}_{i}(\cdot)\}_{i\in\mathcal{G}_{g}}$ at the level $\alpha_{\mathrm{GBC}}$. The corresponding threshold for the $g$th group is given by
$T_{g}=\sup\left\{0<t\leq T_{g,\mathrm{up}}\colon\frac{1+\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq t\}}\leq\alpha_{\mathrm{GBC}}\right\},$ (19)
where $T_{g,\mathrm{up}}$ satisfies $T_{g,\mathrm{up}}<\min_{i\in\mathcal{G}_{g}}\hat{\varphi}_{i}(0.5)$. We define the e-value for all $i\in\mathcal{G}_{g}$ as
@@ 4.3 Structure adaptive multiple testing
where $w_{i}>0$ represents the e-value weight for hypothesis $i$ in group $G_{g}$. Finally, we aggregate all e-values from each group and implement the e-BH procedure. A detailed description of our procedure is given in Algorithm 3.
p-values $p_{1},\dots,p_{n}$, covariates $x_{1},\dots,x_{n}$, group indices $\mathcal{G}_{1},\dots,\mathcal{G}_{G}$, significance levels $\alpha_{\mathrm{GBC}}$ and $\alpha_{\mathrm{eBH}}$
for $g=1,\dots,G$ do
     Compute the estimate of the rejection rule $\hat{\varphi}_{i}(\cdot)$ for $i\in\mathcal{G}_{g}$ using $\mathbf{p}_{-g}$ and $\mathbf{x}_{-g}$.
     Calculate the threshold $T_{g}$ using (19).
     for $i\in\mathcal{G}_{g}$ do
         Calculate the e-value $e_{i}$ using (20).
     end for
end for
Assemble the e-values from all groups.
Run the e-BH procedure utilizing the assembled e-values at the level $\alpha_{\mathrm{eBH}}$.
The indices of rejected hypotheses.
4.3.1 Choice of weights and FDR control
To ensure that the FDR is controlled at the desired level, it is crucial to verify that the e-values defined in equation (20) satisfy Condition (3). We shall show that under certain conditions on the weights, the e-values defined by (20) satisfy (3), and as a result, the corresponding e-BH procedure controls the FDR at the desired level. Before stating the main theorem, let us first introduce some notation. Define $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$, $\mathbf{p}_{g}=\{p_{i}\}_{i\in\mathcal{G}_{g}}$, and $\mathbf{p}_{g,i}$ as the collection of p-values obtained by replacing $p_{i}$ with $\tilde{p}_{i}$ in $\mathbf{p}_{g}$ for $i\in\mathcal{G}_{g}$. Also, let $\mathbf{p}_{-g}=\{p_{i}\}_{i=1}^{n}\setminus\mathbf{p}_{g}$. Due to cross-fitting, the estimated function $\hat{\varphi}_{i}(\cdot)$ for $i\in\mathcal{G}_{g}$ only depends on $\mathbf{p}_{-g}$. Moreover, given the fitted functions $\hat{\varphi}_{i}$ for $i\in\mathcal{G}_{g}$, the threshold $T_{g}$ defined in (19) can be treated as a function of $\mathbf{p}_{g}$. Let $T_{g,j}=T_{g}(\mathbf{p}_{g,j};\{\hat{\varphi}_{l}\}_{l\in\mathcal{G}_{g}})$. We impose the following assumptions.
@@ 4.3.1 Choice of weights and FDR control
Let $\{(p_{i},x_{i})\}$ for $1\leq i\leq n$ be the p-value and covariate pairs.
  1. (A)
    The null pairs $\{(p_{i},x_{i})\}_{i\in\mathcal{H}_{0}}$ are mutually independent.
  2. (B)
    The null pairs $\{(p_{i},x_{i})\}_{i\in\mathcal{H}_{0}}$ are independent of the alternative pairs $\{(p_{i},x_{i})\}_{i\notin\mathcal{H}_{0}}$.
  3. (C)
    For $i\in\mathcal{H}_{0}$, $p_{i}$ is independent of $x_{i}$ and satisfies Condition (8).
@@
Assumption 1 concerns the dependence of the null pairs, which is standard in the literature; see, e.g., Assumption 1 of Ignatiadis and Huber (2021) and Zhao and Zhou (2023). Assumption 2 implies that $P(\varphi_{i}(p_{i},x_{i};\beta)\leq b)\leq P(\varphi_{i}(1-p_{i},x_{i};\beta)\leq b)$ for all $\varphi_{i}(0,x_{i};\beta)\leq b\leq\varphi_{i}(0.5,x_{i};\beta)$, which will be used in our proof. We describe a concrete choice of $\varphi_{i}$ in Section 4.3.2 below.
@@
Suppose Assumptions 1 and 2 hold. If the weights $\{w_{i}\}$ are independent of the p-values and covariate information and satisfy (17), then the e-values defined in (20) fulfills Condition (3).
A naive choice is to set $w_{i}=1$ for all $i=1,2,\cdots,n$, which satisfy (17). However, this choice of weights often leads to low statistical power in simulations. To improve efficiency, we propose a data-dependent approach for constructing the weights. Given the group index $g$, for $g^{\prime}\neq g$, $i\in\mathcal{G}_{g}$ and $j\in\mathcal{G}_{g^{\prime}}$, let $\hat{\varphi}_{j}^{g,i,p}$ be the cross-fitted function obtained by replacing $\mathbf{p}_{g}$ with $\mathbf{p}_{g,i}(p)$, where $\mathbf{p}_{g,i}(p)$ is the collection of p-values with $p_{i}$ replaced by $p$. Define $T_{g^{\prime},j}^{g,i,p}=T_{g^{\prime}}(\mathbf{p}_{g^{\prime},j};\{\hat{\varphi}_{l}^{g,i,p}\}_{l\in\mathcal{G}_{g^{\prime}}})$. For $i\in\mathcal{G}_{g}$ with $1\leq g\leq G$, we propose the following e-value weights:
@@
In the case of $G=2$, we have
for $i\in\mathcal{G}_{1}$ and
for $i\in\mathcal{G}_{2}$. The construction of $w_{i}$ involves taking supremum over $p$, which is crucial for the proof to go through. On the one hand, it renders $w_{i}$ independent of $p_{i}$, which is a useful fact in the proof. On the other hand, it makes the weight sufficiently small in the sense that we can upper bound $w_{i}$ with the term $\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}^{2,i,p}(1-p_{j})\leq T_{1,j}^{2,i,p}\}$ in its denominator replaced by $\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}$, which is another fact used in our argument.
@@
4.3.2 Simulation studies

### v2

@@ 5 Structure-Adaptive Multiple Testing
$T_{g}=\sup\left\{0<t\leq T_{g,\mathrm{up}}\colon\frac{1+\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq t\}}\leq\alpha_{\mathrm{FBC}}\right\},$ (9)
where $T_{g,\mathrm{up}}<\min_{i\in\mathcal{G}_{g}}\hat{\varphi}_{i}(0.5)$. We define the e-value for all $i\in\mathcal{G}_{g}$ as
@@
where $w_{i}>0$ represents the e-value weight for hypothesis $i$ in group $G_{g}$. Finally, we aggregate all e-values from each group and implement the e-BH procedure. A detailed description of our procedure is given in Algorithm G.1 in Supplement G.1.
5.1 Weights and FDR Control
To ensure that the FDR is controlled at the desired level, it is crucial to verify that the e-values defined in equation (10) satisfy Condition (2). We shall show that under certain conditions on the weights, the e-values defined by (10) satisfy (2), and as a result, the corresponding e-BH procedure controls the FDR at the desired level. Before stating the main theorem, let us first introduce some notations. Define $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$, $\mathbf{p}_{g}=\{p_{i}\}_{i\in\mathcal{G}_{g}}$, and $\mathbf{p}_{g,i}$ as the collection of p-values obtained by replacing $p_{i}$ with $\tilde{p}_{i}$ in $\mathbf{p}_{g}$ for $i\in\mathcal{G}_{g}$. Also, let $\mathbf{p}_{-g}=\{p_{i}\}_{i=1}^{n}\setminus\mathbf{p}_{g}$. Due to cross-fitting, the estimated function $\hat{\varphi}_{i}(\cdot)$ for $i\in\mathcal{G}_{g}$ only depends on $\mathbf{p}_{-g}$. Moreover, given the fitted functions $\hat{\varphi}_{i}$ for $i\in\mathcal{G}_{g}$, the threshold $T_{g}$ defined in (9) can be treated as a function of $\mathbf{p}_{g}$. Let $T_{g,j}=T_{g}(\mathbf{p}_{g,j};\{\hat{\varphi}_{l}\}_{l\in\mathcal{G}_{g}})$. We impose the following assumptions.
@@ 5.1 Weights and FDR Control
Let $\{(p_{i},x_{i})\}$ for $1\leq i\leq n$ be the p-value and covariate pairs. (A) The null pairs $\{(p_{i},x_{i})\}_{i\in\mathcal{H}_{0}}$ are mutually independent. (B) The null pairs $\{(p_{i},x_{i})\}_{i\in\mathcal{H}_{0}}$ are independent of the alternative pairs $\{(p_{i},x_{i})\}_{i\notin\mathcal{H}_{0}}$. (C) For $i\in\mathcal{H}_{0}$, $p_{i}$ is independent of $x_{i}$ and satisfies Condition (6).
@@
Assumption 1 concerns the dependence of the null pairs, which is standard in the literature; see, e.g., Assumption 1 of Ignatiadis and Huber (2021) and Zhao and Zhou (2024). Assumption 2 implies that $P(\varphi_{i}(p_{i},x_{i};\beta)\leq b)\leq P(\varphi_{i}(1-p_{i},x_{i};\beta)\leq b)$ for all $\varphi_{i}(0,x_{i};\beta)\leq b\leq\varphi_{i}(0.5,x_{i};\beta)$, which will be used in our proof. We shall describe a concrete choice of $\varphi_{i}$ in Section 5.2 below.
@@
Suppose Assumptions 1 and 2 hold. If the weights $\{w_{i}\}$ are independent of the p-values and covariate information and satisfy
$\sum_{g=1}^{G}n_{g}\max_{i\in\mathcal{G}_{g}}w_{i}\leq n,$ (11)
then the e-values defined in (10) fulfill Condition (2).
A naive choice is to set $w_{i}=1$ for all $i=1,2,\dots,n$, which satisfies (11). However, this choice of weights often leads to low statistical power in simulations. To improve efficiency, we propose a data-dependent approach for constructing the weights. Given the group index $g$, for $g^{\prime}\neq g$, $i\in\mathcal{G}_{g}$ and $j\in\mathcal{G}_{g^{\prime}}$, let $\hat{\varphi}_{j}^{g,i,p}$ be the cross-fitted function obtained by replacing $\mathbf{p}_{g}$ with $\mathbf{p}_{g,i}(p)$, where $\mathbf{p}_{g,i}(p)$ is the collection of p-values with $p_{i}$ replaced by $p$. Define $T_{g^{\prime},j}^{g,i,p}=T_{g^{\prime}}(\mathbf{p}_{g^{\prime},j};\{\hat{\varphi}_{l}^{g,i,p}\}_{l\in\mathcal{G}_{g^{\prime}}})$. For $i\in\mathcal{G}_{g}$ with $1\leq g\leq G$, we propose the following e-value weights:
@@
The construction of $w_{i}$ involves taking the supremum over $p$, which is crucial for the proof to go through. On the one hand, it renders $w_{i}$ independent of $p_{i}$, which is a useful fact in the proof. On the other hand, it makes the weight sufficiently small in the sense that we can upper bound $w_{i}$ with the term $\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}^{g,i,p}(1-p_{j})\leq T_{g^{\prime},j}^{g,i,p}\}$ in its denominator replaced by $\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g^{\prime},j}\}$, which is another fact used in our argument.
@@
The $\tau$-censored weighted BH procedure proposed by Zhao and Zhou (2024) is a variant of the weighted BH procedure that employs a leave-one-out strategy to construct weights. A detailed comparison between our method and theirs is provided in Supplement G.1.
5.2 Simulation Studies